OpenStudy (anonymous):

Product A is twice as expensive as product B. A company buys three times as much of B as it does of A and spends $52 in purchasing an order of these two products . This order includes 20 kg of product A. What is the cost per kg of each product? (please show me working)

OpenStudy (anonymous):

Let the price of product A be x so the price of product B is 2x Let the amount of product A be y so the amount of product B is 3y xy + 6xy = 52

OpenStudy (anonymous):

so product A weighs 20/7 kg and product B weighs 120/7 kg

OpenStudy (anonymous):

the question asks how much they cost

OpenStudy (anonymous):

52/7 divided by 20/7 = $52/20 per kg 312/7 divided by 120/7 = $312/120 per kg

OpenStudy (anonymous):

I may have a mistake

OpenStudy (anonymous):

yes cause those two are somehow wrong

OpenStudy (anonymous):

wait

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

y + 3y = 20 kg y = 5 kg Product A is 5 kg Product B is 15 kg

OpenStudy (anonymous):

52/7 divided by 5 312/7 divided by 15

OpenStudy (anonymous):

anyway ill just give this question up

OpenStudy (anonymous):

Product A and B cost $1.04 and $0.52 per kg respectively.

OpenStudy (anonymous):

wow thanks a lot mate